OpenStudy (jenniferjuice):

What are the solutions of x2 + 8x − 20 = 0? Show all work for full credit

OpenStudy (anonymous):

Do you know how to factor?

OpenStudy (anonymous):

Do you know how to complete the square?

OpenStudy (anonymous):

Do you know the quadratic formula?

OpenStudy (anonymous):

x^2+8x-20=0 (x+10)(x-2)

OpenStudy (jenniferjuice):

okay...so

OpenStudy (anonymous):

If either one is \(0\) then the other is \(0\).

OpenStudy (anonymous):

So you have two equations: \[ x+10=0,\quad x-2=0 \]

OpenStudy (jenniferjuice):

okay ..

OpenStudy (jenniferjuice):

would i be answering them with x=10 and x=2 ?

OpenStudy (anonymous):

10+10 = 0?

OpenStudy (anonymous):

You have committed a crime against math.

OpenStudy (jenniferjuice):

alright alright jeez...

OpenStudy (anonymous):

Subtract 10 from both sides.

OpenStudy (anonymous):

wio what do you mean from 10+10=0 ?!!

OpenStudy (anonymous):

She said \(x=10\). We had \(x+10=0\)